RhondaSommer (rhondasommer):

For each problem: Identify the contradictory statement. Explain why the statement you chose contradicts each of the other two. Describe how the contradictory statement could be changed to make it agree with the other two statements.

RhondaSommer (rhondasommer):



RhondaSommer (rhondasommer):

@marziman

OpenStudy (marziman):

The solution I got was 0 = 0 after subtracting 6 from both sides

OpenStudy (marziman):

That's for statement 1 FYI

OpenStudy (marziman):

I can't solve for the other statements, sorry :(

RhondaSommer (rhondasommer):

b-b-b- MARZIIIIII :(

OpenStudy (marziman):

xD

OpenStudy (marziman):

Do you have Kik by the way?

zepdrix (zepdrix):

Statement 1 says that if you apply polynomial long division,\[\large\rm \frac{x^3-7x+6}{x+4}=x^2-4x+9+\frac{\color{indianred}{-30}}{x+4}\]the remainder is -30.

zepdrix (zepdrix):

Statement 2 says that -4 is a root or zero of the polynomial, \(\large\rm f(-4)=0\)

zepdrix (zepdrix):

Statement 3 says that f(-4) gives us a remainder of -30 by the remainder theorem.

OpenStudy (marziman):

What exactly is it asking? I assumed that we were to solve for each individually. Which is why I solved statement 1 fully.

zepdrix (zepdrix):

It's asking "which statement is lying?"

OpenStudy (marziman):

I assumed that in the beginning but I was unsure. Thank you for the clarification.

OpenStudy (marziman):

Well we know Statement 1 is true.

RhondaSommer (rhondasommer):

well... I guess im confused.

zepdrix (zepdrix):

Still confused? :o Statement 1 is saying that f(x) has a remainder of -30 when x+4 is divided out. Statement 2 is saying that f(x) has a zero at x=-4, no remainder. Statement 3 is saying that f(x) has a remainder of -30 at x=-4. One of those statements ... is not like the others :O

RhondaSommer (rhondasommer):

I think its A... but im really unsure now..

RhondaSommer (rhondasommer):

@zepdrix

zepdrix (zepdrix):

Oh this still -_- ummm

RhondaSommer (rhondasommer):

im sorry :(

zepdrix (zepdrix):

You really don't understand what I'm saying...? The first one says something about a remainder... The third one says something about a remainder... the second one....... does not

RhondaSommer (rhondasommer):

OHHHHHHHHHHHHHHHHHHHHHH I C I C lmao ur smart. could you help me with similar ones?

RhondaSommer (rhondasommer):

oh but first. how could the second one be changed to make it similar

RhondaSommer (rhondasommer):

@3mar

RhondaSommer (rhondasommer):

@steve816

OpenStudy (3mar):

I am sorry. I was not here! If you still need help, I am at your service!

OpenStudy (3mar):

With @zepdrix, you are in good hands @RhondaSommer!